OpenStudy (just_one_last_goodbye):

Solve the system of equations. 2x − 5y = 3 x − 3y = 1

OpenStudy (just_one_last_goodbye):

A. (4, 1) B. (5, 2) C. (7, 2) D. (9, 3)

OpenStudy (anonymous):

Well we have a system of equations here. Do you know how to do the elimination method?

OpenStudy (just_one_last_goodbye):

Process of elimination?

OpenStudy (mertsj):

1. Multiply the bottom equation by -2 2. Add the result to the first equation. 3. Solve the resulting equation for y

OpenStudy (anonymous):

well lol i was gonna say that ^

OpenStudy (mertsj):

4. Substitute that value of y into either equation and find x.

pooja195 (pooja195):

Solve the system of equations. 2x − 5y = 3 -2 (x − 3y = 1) Distribute the -2 \(\huge\color{red}{-2 (x − 3y = 1)}\)

OpenStudy (anonymous):



OpenStudy (just_one_last_goodbye):

-2(-3=y(x) + x)

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin5.htm this will help you understand elimination method

OpenStudy (just_one_last_goodbye):

it's A isn't it?

OpenStudy (anonymous):

yes. but you do u understand elimination

OpenStudy (just_one_last_goodbye):

yes sir thank you so much

OpenStudy (anonymous):

NP!!! ;)

OpenStudy (just_one_last_goodbye):

oh sorry forgot to give u a medal xD